In history class, the girl to boy ratio is 9 to 8.
If there are a total of 68 students, how many girls are there?

Found 2 solutions by ewatrrr, ptaylor:
Answer by ewatrrr(24785)

 
Hi
girl to boy ratio is 9 to 8
Let x represent the constant of proportionality
9x + 8x = 68
17x = 68
x = 4
......
9•4 = 36 Girls
And...checking
Girls + boys
36 + 32 = 68 total

Answer by ptaylor(2198)

Let x=number of girls
Girl to boy ratio is 9 to 8 means that for every (9+8=17) students 9/17 are girls. So
(9/17)*68=36--number of girls
Another way:
9:8=x:(68-x) or
(9/8)=x/(68-x) Cross-multiply
9(68-x)=8x
612-9x=8x
17x=612
x=36--number of girls
Did this help?--ptaylor